# MetaQAP - A Meta-Learning Approach for Quality-Aware Pretraining in Image Quality Assessment

### v1

@@ Abstract
meta-learning, quality-aware pre-training, image quality assessment, authentic distortions, quality-aware loss function.
@@ 3 Proposed Approach
Image Quality Assessment (IQA) is vital for applications like benchmarking, video streaming, and encoding. Addressing IQA’s challenges involves leveraging natural scene statistics, deep learning, or hybrid approaches, categorized as full-reference or no-reference based on reference availability. While full-reference methods are widely adopted, no-reference algorithms often face generalization issues, performing inconsistently across diverse datasets. This study aims to enhance the generalizability of no-reference IQA through a novel approach validated on authentically distorted images and cross-dataset evaluations.
@@
To systematically address IQA, we adapted an established framework known for structuring machine learning projects. Although methodologies like KDD and SEMMA are common, we employed CRISP-DM (Cross Industry Standard Process for Data Mining) [41] due to its versatility and iterative nature.
Our tailored CRISP-DM adaptation for IQA focuses on iterative refinement across key stages: data understanding, preparation, modeling, evaluation, and deployment. This iterative design ensures continuous improvement and robust generalizability.
@@
IQA is a challenging task due to the subjective nature of human perception and the diversity of distortions in real-world images. The primary objective of this study is to develop a no-reference IQA model that accurately predicts perceptual quality scores, aligning them with subjective human ratings (MOS).Training involves annotated datasets with MOS ratings, while validation employs holdout or independently collected datasets. Model performance is evaluated using metrics such as Pearson Linear Correlation Coefficient (PLCC) and Spearman Rank Order Correlation Coefficient (SROCC).
Furthermore, this study emphasizes the importance of addressing existing gaps in no-reference IQA, including poor cross-dataset performance and sensitivity to authentic distortions. By incorporating innovative techniques such as quality-aware pre-training, tailored loss functions, and meta-learning, the proposed model aims to achieve both higher predictive accuracy and broader applicability in diverse real-world scenarios.
@@
The effectiveness of IQA models depends on the diversity and authenticity of the training datasets. This study leverages three benchmark datasets: LiveCD [40], KonIQ-10K [13], and BIQ2021 [2], featuring authentically distorted images to ensure robust model development and evaluation.
@@
Data preparation is the cornerstone of any successful machine learning pipeline, especially for deep learning models like CNNs. Data preprocessing enhances model performance and ensures better generalization across datasets. In this work, CNNs are employed to automatically learn distinguishing features from labeled images, with a significant difference from traditional natural scene statistics-based methods. Inspired by prior research [28, 29, 35, 22, 26], the study has strategically implement data preparation techniques to maximize the performance of proposed quality assessment models. Below are the critical steps that ensure the robustness and reliability of the proposed approach:
Image Resizing and Normalization
While early CNN architectures required input images of fixed dimensions, modern designs incorporate adaptive pooling and fully convolutional layers, allowing them to handle variable image sizes. However, resizing or cropping remains a common preprocessing step to ensure consistency in feature extraction. Resizing may alter perceptual quality, making it less reliable for IQA. Instead, we utilize random cropping, which extracts patches of the desired size. This method acts as a natural regularizer by exposing the model to different regions of the same image during training. Consequently, the model learns to assess perceptual quality independent of content-specific details.
The images in the datasets are 24-bit color images with three layers (red, green, and blue), each with pixel values ranging from 0 to 255. To expedite model convergence and improve numerical stability, these pixel values are normalized by dividing them by 255, bringing them into the range of 0 to 1.
@@
The Mean Opinion Score (MOS), derived from subjective ratings, varies across datasets due to differences in rating scales. To address this and enable consistent cross-dataset evaluation, The MOS score is normalized using min-max scaling:
$MOS_{scaled}=\frac{MOS_{max}-MOS_{min}}{MOS-MOS_{min}}$ (1)
This transformation standardizes the MOS values into a uniform range of 0 to 1, where 0 represents the lowest perceptual quality and 1 represents the highest. This ensures smooth convergence during training and facilitates seamless integration of multiple datasets.
@@
In IQA, traditional augmentation techniques like scaling or contrast adjustments—commonly used in classification tasks—are unsuitable because they may alter the perceptual quality of images. Instead, this study employs carefully selected augmentation methods, including translation, rotation, and random cropping, which preserve perceptual quality while introducing meaningful variability. For instance, random cropping ensures the model encounters different regions of the same image during training, promoting a focus on perceptual quality rather than content-specific details.
@@
For this study, we use 80% of the data for training and 20% for validation, following standard practices in the field.
Given the computational complexity of our approach and trends in the literature, holdout validation is adopted, where 80% of the data is used for training and 20% for testing. Furthermore, in cross-dataset evaluations, the model is trained on 100% of one dataset and tested on 100% of another, ensuring maximum learning from the training data while testing the model’s generalization across datasets.
@@
Sr. Distortion Model Description
1 Color quantization Convert to indexed image using minimum variance quantization and dithering.
2 JPEG2000 Applies standard JPEG2000 compression.
3 Impulse noise Adds salt and pepper noise to the RGB image.
4 Color block Inserts homogeneous random colored blocks at random locations in the image.
5 Color shift Blend green channel randomly into the normalized gradient magnitude of the image.
6 Color diffusion Gaussian blurs the color channels (a and b) in the Lab color space.
7 Multiplicative noise Introduces speckle noise to the RGB image.
8 Denoise Adds Gaussian noise and perform denoising using DuCNN.
9 Gaussian blur Applies a filter with a variable Gaussian kernel to blur the image.
10 High sharpen Over-sharpens the image using unsharp masking.
11 Darken Similar to brighten but decreases other luminance values to darken the image.
12 Contrast change Non-linearly change the color contrast of the image in the form of S-curve.
13 HSV Color saturation Multiplies the saturation channel in the HSV color space by a factor.
14 Jitter Randomly scatter image data with random small offsets using bicubic interpolation.
15 Lens blur Uses a filter with a circular kernel to simulate lens blur.
16 JPEG Uses standard JPEG compression.
17 Mean shift Adds a constant value to all pixels and truncates to the original value range.
18 Brighten Non-linearly adjusts the luminance channel to increase brightness.
19 Lab Color saturation Multiplies the color channels in the Lab color space by a factor.
20 Non-eccentricity patch Randomly offsets small patches in the image to nearby locations.
21 Pixelate Downscale the image & perform upsampling using nearest-neighbor interpolation.
22 Quantization Quantizes image values using N thresholds obtained with Otsu’s method.
23 Motion blur Applies a filter with a line kernel to create motion blur effects.
24 White noise in YCbCr Adds Gaussian noise to the luminance & chrominance channels of the YCbCr image.
25 White noise Adds Gaussian white noise to the RGB image.
@@
The differentiable SROCC component is expressed as $1-SROCC$ to frame it as a loss term (i.e., minimizing it is desirable). The coefficients $\lambda_{1}$ and $\lambda_{2}$ control the relative importance of the MSE and SROCC components, respectively.
@@
Sr. Network Parameters Memory Depth MACs KonIQ-10K LiveCD BIQ2021
(Millions) (MB) (Layers) (GMACs) PLCC SROCC PLCC SROCC PLCC SROCC
1 Densenet-201 20.0 77 201 $\sim$4.4 0.9404 0.9268 0.9093 0.8957 0.7465 0.7329
2 Efficientnet-b0 5.3 20 82 $\sim$0.39 0.8509 0.8265 0.8096 0.7852 0.5995 0.5773
3 GoogleNet 7.0 27 22 $\sim$1.5 0.8867 0.8791 0.8165 0.8089 0.6249 0.6005
4 Inception-V3 23.9 91 48 $\sim$5.7 0.5206 0.6874 0.5528 0.7196 0.6808 0.6732
5 InceptionResNet-V2 55.9 213 164 $\sim$13.2 0.6622 0.7833 0.6796 0.8007 0.3379 0.5047
6 MobileNet-V2 3.5 14 53 $\sim$0.3 0.8626 0.8445 0.8318 0.8137 0.4989 0.6200
7 NasNet_Large 88.9 344 - $\sim$23.8 0.9643 0.9403 0.9717 0.9477 0.6843 0.6662
8 NasNet_Mobile 5.3 20 - $\sim$1.1 0.8195 0.8329 0.7477 0.7611 0.8505 0.8265
9 Resnet18 11.7 45 18 $\sim$1.8 0.7855 0.7896 0.7563 0.7604 0.6067 0.6201
10 Resnet50 25.6 98 50 $\sim$3.9 0.8328 0.8123 0.8483 0.8278 0.6091 0.6132
11 Resnet101 44.6 171 101 $\sim$7.6 0.7119 0.7114 0.7271 0.7266 0.7267 0.7062
12 ShuffleNet 1.4 5.5 50 $\sim$0.15 0.7889 0.7843 0.7380 0.7334 0.5836 0.5831
13 SqueezeNet 1.2 4.7 18 $\sim$0.8 0.7664 0.7583 0.7703 0.7622 0.5481 0.5435
14 Xception 22.9 88 71 $\sim$8.4 0.9279 0.8997 0.9153 0.8871 0.5874 0.5793
@@
10:     Best_score $\leftarrow$ $\infty$- ∞
@@
[…]
@@ 9 ValidationPatience 20
The evaluation of IQA models focuses on their ability to predict quality scores that align with human perceptual judgments. Unlike standard regression problems, where metrics such as MSE and RMSE quantify absolute prediction errors, IQA emphasizes the correlation between predicted scores and ground truth values. To this end, two widely adopted metrics—PLCC and SROCC—are used to evaluate model performance.
@@
PLCC measures the strength and direction of the linear relationship between predicted and observed quality scores. It ranges from $-1$ (perfect negative correlation) to $+1$ (perfect positive correlation), with 00 indicating no linear correlation. This metric evaluates how accurately the model captures linear dependencies in perceptual quality predictions.
@@ 5 Results and Discussion
The proposed model outperformed existing methods across all three benchmark datasets (LiveCD, KonIQ-10K, and BIQ2021) in both PLCC and SROCC metrics, as shown in Table 5. Compared to traditional no-reference IQA methods, such as NIQE, BIQI, and BRISQUE, the proposed model achieved substantial improvements, demonstrating its robustness in predicting quality under authentic distortions. For instance, the proposed model achieved a PLCC of 0.9885 on LiveCD, significantly surpassing the 0.8617 reported by PIQI, a close competitor. These results highlight the effectiveness of integrating pre-training, the quality-aware loss function, and the meta-learner in delivering state-of-the-art performance. These findings demonstrate the robustness and effectiveness of the proposed model across diverse benchmark datasets, highlighting its ability to accurately predict image quality scores under various authentic distortions.
@@
Methods LiveCD KonIQ-10K BIQ2021
PLCC SROCC PLCC SROCC PLCC SROCC
NIQE [46] 0.328 0.299 0.319 0.400 0.301 0.356
GWH-GLBP [47] 0.584 0.559 0.688 0.698 0.644 0.602
BIQI [48] 0.519 0.488 0.718 0.662 0.683 0.564
Robust BRISQUE [46] 0.522 0.484 0.723 0.668 0.664 0.605
PIQE [49] 0.172 0.108 0.208 0.246 0.255 0.213
IL-NIQE [50] 0.487 0.415 0.707 0.447 0.694 0.461
BLIINDS-II [51] 0.473 0.442 0.652 0.575 0.403 0.496
DIIVINE [52] 0.617 0.580 0.463 0.693 0.541 0.617
CurveletQA [53] 0.636 0.621 0.730 0.718 0.698 0.63
OG-IQA [54] 0.545 0.505 0.709 0.635 0.684 0.371
BRISQUE [55] 0.524 0.497 0.574 0.677 0.555 0.603
GM-LOG-BIQA [56] 0.607 0.604 0.589 0.696 0.603 0.617
ENIQA [57] 0.596 0.564 0.761 0.745 0.703 0.634
SSEQ [53] 0.487 0.436 0.637 0.572 0.633 0.528
BMPRI [58] 0.541 0.487 0.705 0.619 0.699 0.494
NBIQA [59] 0.629 0.604 0.771 0.749 0.718 0.642
SGL-IQA [60] 0.704 0.667 0.874 0.794 - 0.710
PIQI [7] 0.862 0.8376 0.815 0.824 0.728 0.745
RAN4IQA [61] - - 0.752 0.763 - -
DB-CNN [62] 0.869 0.851 - 0.875 - -
AIGQA [34] - - 0.868 0.766 - -
DeepEns [22] 0.914 0.899 0.773 0.864 0.791 0.881
NASNet-Large [22] 0.837 0.799 0.790 0.785 0.770 0.750
VRL-IQA [15] 0.941 0.905 0.887 0.884 0.786 0.793
IQDLNet [63] 0.873 0.869 0.921 0.907 - -
TReS [64] 0.877 0.846 0.928 0.915 - -
ARNIQA [65] 0.942 0.937 0.893 0.870 0.845 0.833
UNI-IQA [66] 0.885 0.874 0.937 0.928 0.898 0.907
Q-Align [67] 0.887 0.883 0.935 0.934 - -
MetaQAP 0.988 0.981 0.970 0.966 0.884 0.876

### v2

@@ Abstract
meta-learning, quality-aware pre-training, image quality assessment, authentic distortions, quality-aware loss function
@@ 3 Proposed Approach
This section presents the proposed MetaQAP framework, a no-reference Image Quality Assessment (IQA) method that integrates quality-aware pretraining, meta-learning, and cross-dataset fine-tuning to enhance generalization on authentically distorted images. The approach is designed to systematically bridge the gap between synthetic pretraining and real-world evaluation through an iterative, data-driven learning process.
@@
To ensure a structured and reproducible workflow, this study adapts the CRISP-DM (Cross-Industry Standard Process for Data Mining) framework [41]. While other methodologies such as KDD (Knowledge Discovery in Databases) and SEMMA (Sample, Explore, Modify, Model, Assess) provide systematic approaches to data-driven modeling, CRISP-DM offers greater flexibility and supports iterative refinement—an essential feature for tasks like IQA, where continuous feedback and re-evaluation are required. The proposed adaptation of CRISP-DM for IQA comprises five iterative stages: (1) data understanding, (2) data preparation, (3) modeling, (4) evaluation, and (5) deployment. This structure enables ongoing optimization of feature representations and model generalization.
@@
Image Quality Assessment (IQA) aims to predict human-perceived visual quality without access to a reference image. Existing no-reference (NR) methods often struggle to generalize across datasets, particularly when faced with authentic distortions that differ from synthetic training samples. To address these issues, this study introduces MetaQAP, a NR-IQA framework that combines (i) a quality-aware pretraining stage using synthetically distorted data, (ii) a fine-tuning stage on authentically distorted datasets, and (iii) a meta-learning phase that enhances cross-dataset generalization. The model is trained to align predicted quality scores with human Mean Opinion Scores (MOS), and its performance is measured using Pearson Linear Correlation Coefficient (PLCC) and Spearman Rank Order Correlation Coefficient (SROCC).
@@
The proposed approach is evaluated using three benchmark datasets containing authentically distorted images: LIVE Challenge (LiveCD) [40], KonIQ-10K [13], and BIQ2021 [2]. Each dataset includes Mean Opinion Scores (MOS) and variance values derived from large-scale human studies, providing reliable ground truth for perceptual quality assessment.
@@
Data preparation is crucial for ensuring the robustness and generalizability of deep learning-based IQA models. Unlike traditional natural scene statistics approaches, CNNs automatically learn hierarchical representations of image quality from data. Following prior research [28, 29, 35, 22, 26], this study has strategically implemented preprocessing techniques to optimize model convergence and stability.
Image Resizing, Normalization, and Augmentation
To ensure consistency in feature extraction, images are resized to a fixed dimension while maintaining aspect ratio. Instead of global resizing—which may alter perceptual quality—the model uses random cropping to extract multiple patches from each image. This technique acts as data augmentation by exposing the model to diverse regions of the same image, thereby improving generalization without altering perceptual integrity. Each image is a 24-bit RGB input with pixel values normalized to the [0,1] range by dividing by 255, which stabilizes gradient updates and accelerates training. Additional augmentations, including translation and rotation, further expand the dataset without compromising perceived quality (Figure 4).
@@
The Mean Opinion Score (MOS) varies across datasets due to differing rating scales. To ensure consistent training and evaluation, MOS values are normalized using min–max scaling:
$MOS_{scaled}=\frac{MOS-MOS_{min}}{MOS_{max}-MOS_{min}}$ (1)
This transformation standardizes all MOS values to the [0,1] range, where 0 represents the lowest perceptual quality and 1 the highest. Normalization facilitates smooth convergence during training and enables seamless integration of multiple datasets.
@@
In IQA, traditional augmentation techniques like scaling or contrast adjustments commonly used in classification tasks are unsuitable because they may alter the perceptual quality of images. Instead, this study employs carefully selected augmentation methods, including translation, rotation, and random cropping, which preserve perceptual quality while introducing meaningful variability. For instance, random cropping ensures the model encounters different regions of the same image during training, promoting a focus on perceptual quality rather than content-specific details.
@@
For model training, 80% of each dataset is used for training and 20% for validation. Although cross-validation offers more exhaustive evaluation, it is computationally demanding; therefore, the holdout method is adopted for efficiency. In cross-dataset evaluation, the model is trained on one dataset (e.g. KonIQ-10K) and tested on a distinct dataset (e.g., LIVE Challenge or BIQ2021) to rigorously assess generalization across unseen domains.
@@
Sr. Distortion Model Description 1 Color quantization Convert to indexed image using minimum variance quantization and dithering. 2 JPEG2000 Applies standard JPEG2000 compression. 3 Impulse noise Adds salt and pepper noise to the RGB image. 4 Color block Inserts homogeneous random colored blocks at random locations in the image. 5 Color shift Blend green channel randomly into the normalized gradient magnitude of the image. 6 Color diffusion Gaussian blurs the color channels (a and b) in the Lab color space. 7 Multiplicative noise Introduces speckle noise to the RGB image. 8 Denoise Adds Gaussian noise and perform denoising using DuCNN. 9 Gaussian blur Applies a filter with a variable Gaussian kernel to blur the image. 10 High sharpen Over-sharpens the image using unsharp masking. 11 Darken Similar to brighten but decreases other luminance values to darken the image. 12 Contrast change Non-linearly change the color contrast of the image in the form of S-curve. 13 HSV Color saturation Multiplies the saturation channel in the HSV color space by a factor. 14 Jitter Randomly scatter image data with random small offsets using bicubic interpolation. 15 Lens blur Uses a filter with a circular kernel to simulate lens blur. 16 JPEG Uses standard JPEG compression. 17 Mean shift Adds a constant value to all pixels and truncates to the original value range. 18 Brighten Non-linearly adjusts the luminance channel to increase brightness. 19 Lab Color saturation Multiplies the color channels in the Lab color space by a factor. 20 Non-eccentricity patch Randomly offsets small patches in the image to nearby locations. 21 Pixelate Downscale the image & perform upsampling using nearest-neighbor interpolation. 22 Quantization Quantizes image values using N thresholds obtained with Otsu’s method. 23 Motion blur Applies a filter with a line kernel to create motion blur effects. 24 White noise in YCbCr Adds Gaussian noise to the luminance & chrominance channels of the YCbCr image. 25 White noise Adds Gaussian white noise to the RGB image.
@@
The differentiable SROCC component is expressed as 1SROCC1-SROCC to frame it as a loss term (i.e., minimizing it is desirable). The coefficients $\lambda_{1}$ and $\lambda_{2}$ control the relative importance of the MSE and SROCC components, respectively.
@@
Sr. Network Parameters Memory Depth MACs KonIQ-10K LiveCD BIQ2021 (Millions) (MB) (Layers) (GMACs) PLCC SROCC PLCC SROCC PLCC SROCC 1 Densenet-201 20.0 77 201 $\sim$4.4 0.9404 0.9268 0.9093 0.8957 0.7465 0.7329 2 Efficientnet-b0 5.3 20 82 $\sim$0.39 0.8509 0.8265 0.8096 0.7852 0.5995 0.5773 3 GoogleNet 7.0 27 22 $\sim$1.5 0.8867 0.8791 0.8165 0.8089 0.6249 0.6005 4 Inception-V3 23.9 91 48 $\sim$5.7 0.5206 0.6874 0.5528 0.7196 0.6808 0.6732 5 InceptionResNet-V2 55.9 213 164 $\sim$13.2 0.6622 0.7833 0.6796 0.8007 0.3379 0.5047 6 MobileNet-V2 3.5 14 53 $\sim$0.3 0.8626 0.8445 0.8318 0.8137 0.4989 0.6200 7 NasNet_Large 88.9 344 - $\sim$23.8 0.9643 0.9403 0.9717 0.9477 0.6843 0.6662 8 NasNet_Mobile 5.3 20 - $\sim$1.1 0.8195 0.8329 0.7477 0.7611 0.8505 0.8265 9 Resnet18 11.7 45 18 $\sim$1.8 0.7855 0.7896 0.7563 0.7604 0.6067 0.6201 10 Resnet50 25.6 98 50 $\sim$3.9 0.8328 0.8123 0.8483 0.8278 0.6091 0.6132 11 Resnet101 44.6 171 101 $\sim$7.6 0.7119 0.7114 0.7271 0.7266 0.7267 0.7062 12 ShuffleNet 1.4 5.5 50 $\sim$0.15 0.7889 0.7843 0.7380 0.7334 0.5836 0.5831 13 SqueezeNet 1.2 4.7 18 $\sim$0.8 0.7664 0.7583 0.7703 0.7622 0.5481 0.5435 14 Xception 22.9 88 71 $\sim$8.4 0.9279 0.8997 0.9153 0.8871 0.5874 0.5793
@@
10:  Best_score $\leftarrow$ -$\infty$
@@
13:   New_model.fit(Predictions[{prediction} $\cup$ Selected_predictions], Target)
14:   Score $\leftarrow$ New_model.score(Predictions[{prediction} $\cup$ Selected_predictions], Target)
15:   if $|\text{New\_model.coef}[0]|\geq$ Threshold and Score >> Best_score then
@@ 9 ValidationPatience 20
The evaluation of IQA models focuses on their ability to predict quality scores that align with human perceptual judgments. Unlike standard regression problems, where metrics such as MSE and RMSE quantify absolute prediction errors, IQA emphasizes the correlation between predicted scores and ground truth values. To this end, two widely adopted metrics, PLCC and SROCC, are used to evaluate model performance.
@@
PLCC measures the strength and direction of the linear relationship between predicted and observed quality scores. It ranges from 1-1 (perfect negative correlation) to +1+1 (perfect positive correlation), with 0 indicating no linear correlation. This metric evaluates how accurately the model captures linear dependencies in perceptual quality predictions.
@@ 5 Results and Discussion
The proposed model outperformed existing methods across all three benchmark datasets (LiveCD, KonIQ-10K, and BIQ2021) in both PLCC and SROCC metrics, as shown in Table 5. Compared to traditional no-reference IQA methods, such as NIQE, BIQI, and BRISQUE, the proposed model achieved substantial improvements, demonstrating its robustness in predicting quality under authentic distortions. For instance, the proposed model achieved a PLCC of 0.9885 on LiveCD, significantly surpassing the 0.8617 reported by PIQI, a close competitor. These results highlight the effectiveness of integrating pre-training, the quality-aware loss function, and the meta-learner in delivering state-of-the-art performance. These findings demonstrate the robustness and effectiveness of the proposed model across diverse benchmark datasets, highlighting its ability to accurately predict image quality scores under various authentic distortions. Figure 7 provide the stacked bar-chart depicting the performance of each method in terms of various metrics for an easier comparison.
@@
Methods LiveCD KonIQ-10K BIQ2021 PLCC SROCC PLCC SROCC PLCC SROCC NIQE [46] 0.328 0.299 0.319 0.400 0.301 0.356 GWH-GLBP [47] 0.584 0.559 0.688 0.698 0.644 0.602 BIQI [48] 0.519 0.488 0.718 0.662 0.683 0.564 Robust BRISQUE [46] 0.522 0.484 0.723 0.668 0.664 0.605 PIQE [49] 0.172 0.108 0.208 0.246 0.255 0.213 IL-NIQE [50] 0.487 0.415 0.707 0.447 0.694 0.461 BLIINDS-II [51] 0.473 0.442 0.652 0.575 0.403 0.496 DIIVINE [52] 0.617 0.580 0.463 0.693 0.541 0.617 CurveletQA [53] 0.636 0.621 0.730 0.718 0.698 0.630 OG-IQA [54] 0.545 0.505 0.709 0.635 0.684 0.371 BRISQUE [55] 0.524 0.497 0.574 0.677 0.555 0.603 GM-LOG-BIQA [56] 0.607 0.604 0.589 0.696 0.603 0.617 ENIQA [57] 0.596 0.564 0.761 0.745 0.703 0.634 SSEQ [53] 0.487 0.436 0.637 0.572 0.633 0.528 BMPRI [58] 0.541 0.487 0.705 0.619 0.699 0.494 NBIQA [59] 0.629 0.604 0.771 0.749 0.718 0.642 SGL-IQA [60] 0.704 0.667 0.874 0.794 0.710 PIQI [7] 0.862 0.838 0.815 0.824 0.728 0.745 RAN4IQA [61] 0.752 0.763 DB-CNN [62] 0.869 0.851 0.875 AIGQA [34] 0.868 0.766 DeepEns [22] 0.914 0.899 0.773 0.864 0.791 0.881 NASNet-Large [22] 0.837 0.799 0.790 0.785 0.770 0.750 VRL-IQA [15] 0.941 0.905 0.887 0.884 0.786 0.793 IQDLNet [63] 0.873 0.869 0.921 0.907 TReS [64] 0.877 0.846 0.928 0.915 ARNIQA [65] 0.942 0.937 0.893 0.870 0.845 0.833 UNI-IQA [66] 0.885 0.874 0.937 0.928 0.898 0.907 Q-Align [67] 0.887 0.883 0.935 0.934 MetaQAP (Ours) 0.988 0.981 0.970 0.966 0.884 0.876
Figure 7: Comparison of various models using stacked bar-chart to highlight overall performance